# Tidal Love numbers and Green’s functions in black hole space-times

### v1

@@ Abstract
Tidal interactions play a crucial role in deciphering gravitational wave signals emitted by the coalescence of binary systems. They are usually quantified by a set of complex coefficients which include tidal Love numbers, describing the conservative response to an external perturbation. In the static case, these are found to vanish exactly for asymptotically flat black holes in general relativity in four space-time dimensions, and recently they have been generalized to dynamical interactions. In the context of response theory, the retarded Green’s function provides the complete description of the behavior of dynamical systems. In this work we investigate the relation between Love numbers and Green’s functions, and highlight the relevance of radiation reaction effects to their connection. As a special case, we discuss BTZ black holes, where the absence of radiative modes allows us to make a direct link between them.
@@ I Introduction
An enduring analytical framework to characterize tidal effects is a set of complex coefficients. Their real parts, often called tidal Love numbers (TLNs), describe the linear conservative response of self-gravitating bodies Love (1909); their imaginary parts are instead associated to dissipative effects. Initially formulated within the realm of Newtonian gravity, TLNs have undergone further developments to encompass a fully relativistic context Hinderer (2008); Binnington and Poisson (2009); Damour and Nagar (2009). Within this framework, it is widely recognized that the static TLNs of different families of asymptotically flat BHs in four-dimensional space-time, such as Schwarzschild, Kerr and Reissner-Nordstrom BHs, are precisely zero Binnington and Poisson (2009); Damour and Nagar (2009); Damour and Lecian (2009); Pani et al. (2015a, b); Gürlebeck (2015); Porto (2016a); Le Tiec and Casals (2021); Chia (2021); Le Tiec et al. (2021); Charalambous et al. (2021a, b); Ivanov and Zhou (2023a); Charalambous et al. (2022); Katagiri et al. (2023); Ivanov and Zhou (2023b); Berens et al. (2023); Bhatt et al. (2023); Sharma et al. (2024); Rai and Santoni (2024). However, this characteristic is fragile, as it is violated in scenarios involving BH mimickers Pani (2015); Cardoso et al. (2017), in the presence of a cosmological constant Nair et al. (2024) or extended gravitational theories Cardoso et al. (2017, 2018); De Luca et al. (2023a); Barura et al. (2024), in higher dimensions Kol and Smolkin (2012); Cardoso et al. (2019); Hui et al. (2021); Rodriguez et al. (2023); Charalambous and Ivanov (2023); Charalambous (2024a, b), or in non-vacuum environments induced by secular effects such as accretion or superradiant instabilities of ultralight bosonic fields Baumann et al. (2019); Cardoso and Duque (2020); De Luca and Pani (2021); Cardoso et al. (2022); De Luca et al. (2023b).
@@
Among these contributions, one of the most significant is radiation reaction. The concept of radiation reaction in gravity plays a crucial role in the study of coalescing binary systems (see Refs. Detweiler (2005); Poisson et al. (2011); Barack and Pound (2019) for reviews). There, the energy and momentum carried away by the emitted gravitational waves result in the binary constituents experiencing a reaction force, which influences their motion and changes their trajectory and energy. This effect is completely analogous to the radiation reaction in classical electrodynamics, leading to the familiar Abraham-Lorentz force. The most striking application concerns extreme mass-ratio inspirals, describing the evolution of a binary system with a large hierarchy in masses between its constituents. At leading order in the mass ratio, one has a point-like particle moving in a geodesic orbit around the heavier BH, while at subsequent orders the interaction of the particle with its own gravitational perturbation gives rise to an effective “self-force”, which changes the evolution of the orbit.
@@
The outline of the paper is the following. In Sec. II we discuss the retarded Green’s function in linear response theory. Its connection to TLNs, as well as a review on the topic, is performed in Sec. III for Newtonian gravity. In Sec. IV we overview the formalism to extract TLNs in general relativity, which we then apply to Schwarzschild BHs. In Sec. V we point out a connection between tidal response and tail effects, discussing the special case of BTZ BHs in Sec. VI. The conclusions are left to Sec. VII. In this paper we use natural units $\hbar=c=1$, and mostly-positive signature.
@@ II Retarded Green’s function in linear response theory
where $x$ is the tortoise coordinate for the space-time of interest, $\tilde{R}(x)$ is the (rescaled) radial part of the field profile, and $V$ is an effective potential encoding the information of the background geometry. Examples of wavelike equations of this form are those for odd (Regge-Wheeler) Regge and Wheeler (1957) and even (Zerilli) Zerilli (1970) gravitational perturbations on Schwarzschild BHs, or the Teukolsky equations for perturbations on a Kerr geometry Teukolsky (1973). Later we will explicitly consider equations of this type for Kerr and BTZ BHs.
@@
  • Firstly, there is the high-frequency contribution, which consists of radiation that propagates directly from the source to the observer, and is associated with the large-frequency arc that closes the contour in the complex $\omega$ plane. This contribution reduces to the free-space Green’s function in the limit of vanishing BH mass Leaver (1986).
  • The second contribution comes from quasinormal modes (QNMs). These are singularities in the lower half complex plane (as expected from a retarded Green’s function), which occur at frequencies where the Wronskian vanishes, i.e., where the two solutions become linearly dependent. They describe the interaction of the signal from the source with the curved BH geometry, and the fraction of the ingoing initial perturbation that is eventually reflected from the curvature potential near the event horizon back to spatial infinity Berti et al. (2009).
  • Lastly, there is a contribution describing the late-time behavior of the Green’s function, where the dynamics starts to be affected by radiation reaction effects Detweiler (2005); Poisson et al. (2011); Barack and Pound (2019). Among them, tail effects depend on the asymptotic behavior of the effective potential Ching et al. (1995) and account for the wave scattering off the weak Coulomb potential near infinity Leaver (1986). In other words, they account for the propagation of waves on the light-cones of the BH geometry in linear theory, rather than those of Minkowski space-time. For example, for asymptotically flat space-times, the power-law decrease of the potential at spatial infinity gives rise to late-time tails of the retarded Green’s function Price (1972), which are usually associated with the existence of a branch cut along the negative imaginary axis in complex frequency space.
@@
Of these three contributions, the first does not contain any information about the BH, while the other two encode the response of the object to the external source. This is transparent in the asymptotic approximation, in which it is assumed that the space-time is essentially flat around the observer and the source, both located far away from the BH ($x_{2},x_{1}\gg x_{+}$, but keeping the hierarchy $x_{2}\gtrsim x_{1}$). In this regime, the Green’s function reduces to Andersson (1997)
@@ III Tidal response and Green’s functions in Newtonian gravity
Consider a spherically-symmetric spinning body of mass $M$, assumed to be at the origin of a Cartesian coordinate frame, which is subjected to an external tidal gravitational field $U_{\text{\tiny ext}}$ applied adiabatically. Within spherical symmetry, the gravitational field can be expanded in multipole moments as
@@ IV.1 Static TLNs and effective field theory
where the ellipses denote subleading powers of $r_{+}/r$. Following the Newtonian matching procedure, one can make an analogy with the non-relativistic gravitational potential of Eq. (11) to recognize the external tidal field in the growing term $\sim r^{\ell}$, and the BH response in the decaying one $\sim r^{-\ell-1}$. The coefficients $\kappa_{\ell}$ denote the static TLNs defined in Eq. (13). These are found to vanish identically for Kerr geometries Hui et al. (2021), as we will review in Sec. IV.2.111The overlap between the source series and the response contribution in the physical case $\ell\in\mathbb{N}$ Kol and Smolkin (2012); Le Tiec and Casals (2021); Charalambous et al. (2021a) requires us to perform an analytic continuation to the unphysical region $\ell\in\mathbb{R}$ Le Tiec and Casals (2021), where the source and response series do not overlap.
Tidal effects can also be described in the point-particle effective field theory (EFT) formalism Porto (2016b); Goldberger and Rothstein (2006, 2020); Goldberger et al. (2021). This framework is based on the realization that a BH behaves as a point particle at large distances $r\gg r_{+}$, and corrections due to its internal structure are encoded in higher-derivative operators in the effective theory.
@@ IV.2 Dynamical response and asymptotic matching
As discussed in the previous section, the response to tidal sources is captured by a set of coefficients, which describe both the static and dynamical response to external perturbations. The latter is based on frequency dependent perturbers Perry and Rodriguez (2023), and usually appears as the physically-relevant scenario in the context of binary systems during the inspiral phase. Finding the dynamical response has, however, been shown to suffer from a major drawback due to the difficulty of solving the perturbation equations at finite frequency, which has been achieved only in certain limiting regimes.
The standard approach to treat time-dependent tidal excitations of spherically symmetric BHs invokes the technique of matched asymptotic expansions Poisson (2021); Hegade K. R. et al. (2024), based on the separation of scales schematically shown in Fig. 2. This formalism relies on solving the perturbation equations in three distinct space-time regions, and matching the solutions inside regions of common overlap. Before describing these regions, let us stress that this approach is also based on a perturbative series in the parameter $\omega M\ll 1$, which is equivalent to an expansion in the strength of the external tidal source.
@@
  • The radius of curvature, $\mathcal{R}$, of the external space-time in which the body is inserted.
@@
where ${}_{2}F_{1}$ is the hypergeometric function of the $2^{\rm nd}$ kind, $C$ is a constant, and
@@
We can now determine the TLNs associated to the solution (46). Following the asymptotic matching prescription, we can expand the solution of the inner region to large radii in the buffer region, in order to recognize the growing and decaying branches. In practice, this amounts to using the Kummer property Abramowitz and Stegun (1964)
@@
It is instructive to expand this expression for small spin $\chi$ and frequency $\omega$, using the expression for $\xi$ below Eq. (IV.2). The result is Kehagias et al. (2023):
@@
For integer $\ell$, the overall prefactor is finite, hence the first term (real part) vanishes. Matching to (13), this implies that the conservative response of Kerr BHs vanishes in the static limit, $\kappa_{\ell m}=0$. The second term (imaginary part), which encodes the dissipative part $\nu_{\ell m}$, is non-vanishing and proportional to the frame-dragging term $\omega r_{+}^{2}-m\chi$ Kehagias et al. (2023). Finally, the dynamical TLNs for Schwarzschild BHs are found to vanish within this approximation Charalambous et al. (2021a); Poisson (2021), even though $\mathcal{O}(\omega^{2}r_{+}^{2})$ corrections to the approximated equation of motion could affect this result. In fact, in the context of scattering amplitudes, the dynamical TLNs were instead found to be nonvanishing and characterised by a logarithmic running Saketh et al. (2024).
@@ V.1 Radiation reaction and tail effects
Radiation reaction accounts for the contribution of radiative modes of the outer region to the dynamics in the inner region. It can be accounted for in an EFT approach by integrating out the wave zone contributions. Similarly to the extraction of the dynamical response, the theory of the gravitational self-force in curved space-time is based on the formalism of matched asymptotic expansions and singular perturbation theory Kevorkian and Cole (1996). In that context, one must glue the metric of the BH perturbed by the external universe (source) through which it is moving, with the metric of the external universe once perturbed by the BH moving through it Detweiler (2005).
@@
  • The regular field is a smooth vacuum solution, and contains the backscattered waves that arise from propagation within the light cones of the background space-time.
@@
such that $G_{\text{\tiny ret}}=G_{\text{\tiny sing}}+G_{\text{\tiny reg}}$. [^4] In curved space-time, the construction of the singular Green’s function $G_{\text{\tiny sing}}$ demands the addition of a specific solution to the homogeneous equation satisfying characteristic boundary conditions on the null cones emanating from the source location Poisson et al. (2011).
@@
Since tail effects involve backscattering off the slightly curved space-time far from the BH, one can focus on distances $r\gg r_{s}$ (that is, in the outer region). In this region, the above equation simplifies to
@@
This is compatible with the solution obtained in the MST approach Mano and Takasugi (1997); Mano et al. (1996a, b). In the limit $\omega r_{s}\ll 1$, one easily recovers the known Bessel functions of asymptotically flat space-times. [^5] In the context of a scattering scalar wave on a BH background, this solution can be used to relate the tidal response to the reflection amplitude Ivanov and Zhou (2023b); Charalambous (2024b), showing that vanishing TLNs can be interpreted as reflectionless, total transmission modes Hod (2013); Cook and Zalutskiy (2016a, b).
@@
where $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$ denotes the digamma function. At order $(\omega r_{s})^{0}$, one is left with the first two lines, where one recognizes a factorization of the Green’s function into a source (growing terms) and a response (decaying terms), similar to the asymptotic expression shown in Eq. (7). However, the non-zero BH mass brings an additional contribution to the radial dependence, proportional to a logarithmic term. This is in contrast with the instantaneous decomposition of the Green’s function outlined within the EFT approach (see Eq. (41)), which makes the extraction of the tidal response unclear.

### v2

# Tidal Love numbers and Green’s functions in black hole spacetimes

@@ Abstract
Tidal interactions play a crucial role in deciphering gravitational wave signals emitted by the coalescence of binary systems. They are usually quantified by a set of complex coefficients which include tidal Love numbers, describing the conservative response to an external perturbation. In the static case, these are found to vanish exactly for asymptotically flat black holes in general relativity in four spacetime dimensions, and recently they have been generalized to dynamical interactions. In the context of response theory, the retarded Green’s function provides the complete description of the behavior of dynamical systems. In this work we investigate the relation between Love numbers and Green’s functions, and highlight the relevance of radiation reaction effects to their connection. As a special case, we discuss Banados-Teitelboim-Zanelli black holes, where the absence of radiative modes allows us to make a direct link between them.
@@ I Introduction
An enduring analytical framework to characterize tidal effects is a set of complex coefficients. Their real parts, often called tidal Love numbers (TLNs), describe the linear conservative response of self-gravitating bodies Love (1909); their imaginary parts are instead associated to dissipative effects. Initially formulated within the realm of Newtonian gravity, TLNs have undergone further developments to encompass a fully relativistic context Hinderer (2008); Binnington and Poisson (2009); Damour and Nagar (2009). Within this framework, it is widely recognized that the static TLNs of different families of asymptotically flat BHs in four-dimensional spacetime, such as Schwarzschild, Kerr and Reissner-Nordstrom BHs, are precisely zero Binnington and Poisson (2009); Damour and Nagar (2009); Damour and Lecian (2009); Pani et al. (2015a, b); Gürlebeck (2015); Porto (2016a); Le Tiec and Casals (2021); Chia (2021); Le Tiec et al. (2021); Charalambous et al. (2021a, b); Bonelli et al. (2022); Ivanov and Zhou (2023a); Charalambous et al. (2022); Katagiri et al. (2023); Ivanov and Zhou (2023b); Berens et al. (2023); Bhatt et al. (2023); Sharma et al. (2024); Rai and Santoni (2024). However, this characteristic is fragile, as it is violated in scenarios involving BH mimickers Pani (2015); Cardoso et al. (2017), in the presence of a cosmological constant Nair et al. (2024) or extended gravitational theories Cardoso et al. (2017, 2018); De Luca et al. (2023a); Barura et al. (2024), in higher dimensions Kol and Smolkin (2012); Cardoso et al. (2019); Hui et al. (2021); Rodriguez et al. (2023); Charalambous and Ivanov (2023); Charalambous (2024a, b), or in nonvacuum environments induced by secular effects such as accretion or superradiant instabilities of ultralight bosonic fields Baumann et al. (2019); Cardoso and Duque (2020); De Luca and Pani (2021); Cardoso et al. (2022); De Luca et al. (2023b).
@@
Among these contributions, one of the most significant is radiation reaction. The concept of radiation reaction in gravity plays a crucial role in the study of coalescing binary systems (see Refs. Detweiler (2005); Poisson et al. (2011); Barack and Pound (2019) for reviews). There, the energy and momentum carried away by the emitted gravitational waves result in the binary constituents experiencing a reaction force, which influences their motion and changes their trajectory and energy. This effect is completely analogous to the radiation reaction in classical electrodynamics, leading to the familiar Abraham-Lorentz force. The most striking application concerns extreme mass-ratio inspirals, describing the evolution of a binary system with a large hierarchy in masses between its constituents. At leading order in the mass ratio, one has a pointlike particle moving in a geodesic orbit around the heavier BH, while at subsequent orders the interaction of the particle with its own gravitational perturbation gives rise to an effective “self-force”, which changes the evolution of the orbit.
@@
The outline of the paper is the following. In Sec. II we discuss the retarded Green’s function in linear response theory. Its connection to TLNs, as well as a review on the topic, is performed in Sec. III for Newtonian gravity. In Sec. IV we overview the formalism to extract TLNs in general relativity, which we then apply to Schwarzschild BHs. In Sec. V we point out a connection between tidal response and tail effects, discussing the special case of Banados-Teitelboim-Zanelli (BTZ) BHs in Sec. VI. The conclusions are left to Sec. VII. In this paper we use natural units $\hbar=c=1$, and mostly positive signature.
@@ II Retarded Green’s function in linear response theory
where $x$ is the tortoise coordinate for the spacetime of interest, $\tilde{R}(x)$ is the (rescaled) radial part of the field profile, and $V$ is an effective potential encoding the information of the background geometry. Examples of wavelike equations of this form are those for odd (Regge-Wheeler) Regge and Wheeler (1957) and even (Zerilli) Zerilli (1970) gravitational perturbations on Schwarzschild BHs, or the Teukolsky equations for perturbations on a Kerr geometry Teukolsky (1973). Later we will explicitly consider equations of this type for Kerr and BTZ BHs.
@@
  • First, there is the high-frequency contribution, which consists of radiation that propagates directly from the source to the observer, and is associated with the large-frequency arc that closes the contour in the complex $\omega$ plane. This contribution reduces to the free-space Green’s function in the limit of vanishing BH mass Leaver (1986).
  • The second contribution comes from quasinormal modes (QNMs). These are singularities in the lower half complex plane (as expected from a retarded Green’s function), which occur at frequencies where the Wronskian vanishes, i.e., where the two solutions become linearly dependent. They describe the interaction of the signal from the source with the curved BH geometry, and the fraction of the ingoing initial perturbation that is eventually reflected from the curvature potential near the event horizon back to spatial infinity Leaver (1986); Berti et al. (2009); Hui et al. (2019).
  • Lastly, there is a contribution describing the late-time behavior of the Green’s function, where the dynamics starts to be affected by radiation reaction effects Detweiler (2005); Poisson et al. (2011); Barack and Pound (2019). Among them, tail effects depend on the asymptotic behavior of the effective potential Ching et al. (1995) and account for the wave scattering off the weak Coulomb potential near infinity Leaver (1986). In other words, they account for the propagation of waves on the light-cones of the BH geometry in linear theory, rather than those of Minkowski spacetime. For example, for asymptotically flat spacetimes, the power-law decrease of the potential at spatial infinity gives rise to late-time tails of the retarded Green’s function Price (1972), which are usually associated with the existence of a branch cut along the negative imaginary axis in complex frequency space.
@@
Of these three contributions, the first does not contain any information about the BH, while the other two encode the response of the object to the external source. This is transparent in the asymptotic approximation, in which it is assumed that the spacetime is essentially flat around the observer and the source, both located far away from the BH ($x_{2},x_{1}\gg x_{+}$, but keeping the hierarchy $x_{2}\gtrsim x_{1}$). In this regime, the Green’s function reduces to Andersson (1997)
@@ III Tidal response and Green’s functions in Newtonian gravity
Consider a spherically symmetric spinning body of mass $M$, assumed to be at the origin of a Cartesian coordinate frame, which is subjected to an external tidal gravitational field $U_{\text{\tiny ext}}$ applied adiabatically. Within spherical symmetry, the gravitational field can be expanded in multipole moments as
@@ IV.1 Static TLNs and effective field theory
where the ellipses denote subleading powers of $r_{+}/r$. Following the Newtonian matching procedure, one can make an analogy with the nonrelativistic gravitational potential of Eq. (11) to recognize the external tidal field in the growing term $\sim r^{\ell}$, and the BH response in the decaying one $\sim r^{-\ell-1}$. The coefficients $\kappa_{\ell}$ denote the static TLNs defined in Eq. (13). These are found to vanish identically for Kerr geometries Hui et al. (2021), as we will review in Sec. IV.2.111The overlap between the source series and the response contribution in the physical case $\ell\in\mathbb{N}$ Kol and Smolkin (2012); Le Tiec and Casals (2021); Charalambous et al. (2021a) requires us to perform an analytic continuation to the unphysical region $\ell\in\mathbb{R}$ Le Tiec and Casals (2021), where the source and response series do not overlap.
Tidal effects can also be described in the point-particle effective field theory (EFT) formalism Goldberger and Rothstein (2006); Chakrabarti et al. (2013); Porto (2016b); Goldberger and Rothstein (2020); Goldberger et al. (2021); Creci et al. (2021). This framework is based on the realization that a BH behaves as a point particle at large distances $r\gg r_{+}$, and corrections due to its internal structure are encoded in higher-derivative operators in the effective theory.
@@ IV.2 Dynamical response and asymptotic matching
As discussed in the previous section, the response to tidal sources is captured by a set of coefficients, which describe both the static and dynamical response to external perturbations. The latter is based on frequency dependent perturbers Perry and Rodriguez (2023), and usually appears as the physically relevant scenario in the context of binary systems during the inspiral phase. Finding the dynamical response has, however, been shown to suffer from a major drawback due to the difficulty of solving the perturbation equations at finite frequency, which has been achieved only in certain limiting regimes.
The standard approach to treat time-dependent tidal excitations of spherically symmetric BHs invokes the technique of matched asymptotic expansions Poisson (2021); Hegade K. R. et al. (2024), based on the separation of scales schematically shown in Fig. 2. This formalism relies on solving the perturbation equations in three distinct spacetime regions, and matching the solutions inside regions of common overlap. Before describing these regions, let us stress that this approach is also based on a perturbative series in the parameter $\omega M\ll 1$, which is equivalent to an expansion in the strength of the external tidal source.
@@
  • The radius of curvature, $\mathcal{R}$, of the external spacetime in which the body is inserted.
@@
where ${}_{2}F_{1}$ is the hypergeometric function of the $2{\rm nd}$ kind, $C$ is a constant, and
@@
We can now determine the TLNs associated to the solution (46). Following the asymptotic matching prescription, we can expand the solution of the inner region to large radii in the buffer region, in order to recognize the growing and decaying branches [^4] Recognizing the growing and decaying branches in the field solution relies on the assumption of weak tidal perturbations, which is hypothesized within the asymptotic matching procedure. In the time dependent problem (see footnote 1 for the static case), going to higher orders in frequency may result into a source-response mixing, complicating the extraction of the dynamical TLNs. Similarly, gauge issues, associated to the matching of the field solutions of the inner and outer regions Poisson (2021); Hegade K. R. et al. (2024), could bring further ambiguities.. In practice, this amounts to using the Kummer property Abramowitz and Stegun (1964)
@@
It is instructive to expand this expression for small spin $\chi$ and frequency $\omega$, using the expression for $\xi$ below Eq. (IV.2). The result is Chia (2021); Charalambous et al. (2021a); Kehagias et al. (2023):
@@
For integer $\ell$, the overall prefactor is finite, hence the first term (real part) vanishes. Matching to (13), this implies that the conservative response of Kerr BHs vanishes in the static limit, $\kappa_{\ell m}=0$. The second term (imaginary part), which encodes the dissipative part $\nu_{\ell m}$, is nonvanishing and proportional to the frame-dragging term $\omega r_{+}^{2}-m\chi$ Chia (2021); Charalambous et al. (2021a); Kehagias et al. (2023). Finally, the dynamical TLNs for Schwarzschild BHs are found to vanish within this approximation Charalambous et al. (2021a); Poisson (2021), even though $\mathcal{O}(\omega^{2}r_{+}^{2})$ corrections to the approximated equation of motion could affect this result. In fact, in the context of scattering amplitudes, the dynamical TLNs were instead found to be nonvanishing and characterised by a logarithmic running Saketh et al. (2024).
@@ V.1 Radiation reaction and tail effects
Radiation reaction accounts for the contribution of radiative modes of the outer region to the dynamics in the inner region. It can be accounted for in an EFT approach by integrating out the wave zone contributions. Similarly to the extraction of the dynamical response, the theory of the gravitational self-force in curved spacetime is based on the formalism of matched asymptotic expansions and singular perturbation theory Kevorkian and Cole (1996). In that context, one must glue the metric of the BH perturbed by the external universe (source) through which it is moving, with the metric of the external universe once perturbed by the BH moving through it Detweiler (2005).
@@
  • The regular field is a smooth vacuum solution, and contains the backscattered waves that arise from propagation within the light cones of the background spacetime.
@@
such that $G_{\text{\tiny ret}}=G_{\text{\tiny sing}}+G_{\text{\tiny reg}}$. [^5] In curved spacetime, the construction of the singular Green’s function $G_{\text{\tiny sing}}$ demands the addition of a specific solution to the homogeneous equation satisfying characteristic boundary conditions on the null cones emanating from the source location Poisson et al. (2011).
@@
Since tail effects involve backscattering off the slightly curved spacetime far from the BH, one can focus on distances $r\gg r_{s}$ (that is, in the outer region). In this region, the above equation simplifies to
@@
This is compatible with the solution obtained in the MST approach Mano and Takasugi (1997); Mano et al. (1996a, b). In the limit $\omega r_{s}\ll 1$, one easily recovers the known Bessel functions of asymptotically flat spacetimes. [^6] In the context of a scattering scalar wave on a BH background, this solution can be used to relate the tidal response to the reflection amplitude Ivanov and Zhou (2023b); Charalambous (2024b), showing that vanishing TLNs can be interpreted as reflectionless, total transmission modes Hod (2013); Cook and Zalutskiy (2016a, b).
@@
where $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$ denotes the digamma function. At order $(\omega r_{s})^{0}$, one is left with the first two lines, where one recognizes a factorization of the Green’s function into a source (growing terms) and a response (decaying terms), similar to the asymptotic expression shown in Eq. (7). However, the nonzero BH mass brings an additional contribution to the radial dependence, proportional to a logarithmic term. This is in contrast with the instantaneous decomposition of the Green’s function outlined within the EFT approach (see Eq. (41)), which makes the extraction of the tidal response unclear.